OpenStudy (lgbasallote):

Prove/disprove that \(\sqrt 3\) is irrational

OpenStudy (zugzwang):

Things like this are just begging to be proven by contradiction...

OpenStudy (anonymous):

it's irrational by definition

OpenStudy (zugzwang):

Just go for assuming that it's rational, and then look for a contradiction...

OpenStudy (lgbasallote):

taking the easy way out huh @nphuongsun93

OpenStudy (anonymous):

:D.. *poke google* < http://en.wikipedia.org/wiki/Square_root_of_3#Proof_of_irrationality >

OpenStudy (lgbasallote):

\[\sqrt 3 = \frac ab\] where a/b is in simplest terms \[3 = \frac{a^2}{b^2}\] \[3b^2 = a^2\] then?

OpenStudy (lgbasallote):

@nphuongsun93 don't spoil the steps

OpenStudy (zugzwang):

Well, if a and b are relatively prime, what can you say about a^2 and b^2?

OpenStudy (anonymous):

depends on what you can use easiest way is to assume \[3a^2=b^2\] then by the fundamental theorem of arithmetic, factor \(a^2\) and\ \(b^2\) uniquely into products of primes observe the right hand side has an even number of factors of 3 whereas the left hand side has an odd number of them

OpenStudy (lgbasallote):

i can say that a and b cannot both be even...

OpenStudy (zugzwang):

Their squares must also be relatively prime, right?

OpenStudy (lgbasallote):

by relatively prime you mean...? i'm not really a math lover like you (as you can tell from my profile picture)

OpenStudy (zugzwang):

Relatively prime, their greatest common factor is 1. It's a consequence of a/b being in simplest terms..

OpenStudy (lgbasallote):

so a and b cannot both be even....

OpenStudy (lgbasallote):

same meaning......

OpenStudy (zugzwang):

That follows from them being relatively prime.

OpenStudy (lgbasallote):

as long as they're same...the end justifies the means

OpenStudy (zugzwang):

Well, there you have it, then.

OpenStudy (lgbasallote):

hmm... i think i see the contradiction now

OpenStudy (lgbasallote):

a and b cannot both be even so if b is odd 3b^2 is odd but 3b^2 = a^2 and a^2 is even odd = even <--contradiction if a is odd 3b^2 is even 3b^2 = a^2 and a^2 is odd even = odd same thing

OpenStudy (zugzwang):

uhh... a and b cannot both be even OK if b is odd, then 3b^2 is odd OK 3b^2 = a^2, a^2 is even... No it isn't... you don't know what it is yet.

OpenStudy (lgbasallote):

i just said a and b cannot both be even...so if b is odd, then a is even. if a is even then a^2 is even....

OpenStudy (zugzwang):

If b is odd, then a is even? I think you just proved that if a is odd, then b is odd,

OpenStudy (lgbasallote):

uhh what??

OpenStudy (lgbasallote):

i just said that a and b can't both be even...when did i say they're both odd?

OpenStudy (zugzwang):

You didn't. You didn't say that if b is odd, then a is even.

OpenStudy (lgbasallote):

i said... a and b can't both be even...in other words...either a is even and b is odd...or a is odd and b is even

OpenStudy (lgbasallote):

that was the definition of rational number a/b where a and b are integers and a and b are not both even (because if it's even then it's factorable and not in lowest terms)

OpenStudy (zugzwang):

Look carefully a and b are not both even, in symbols, this is ~(a is even AND b is even) (a is odd OR b is odd) it is not the same as (a is even OR b is even)

OpenStudy (zugzwang):

In other words, it can be the case that both are odd.

OpenStudy (lgbasallote):

yes....i wasn't there yet....i was just proving case 1 and 2

OpenStudy (lgbasallote):

you instantly called my case 1 wrong already

OpenStudy (zugzwang):

In any case, if a and b are relatively prime, then so are a^2 and b^2 but a^2/b^2 = 3, then a^2 = 3b^2 Then a^2/b^2 = 3b^2/b^2... means a^2 and b^2 are not relatively prime, which is a contradiction...

OpenStudy (lgbasallote):

i just said don't spoil the steps.... anyway..you've done it....so how a^2/b^2 = 3b^2/b^2 are not relatively prime?

OpenStudy (zugzwang):

Well, b^2 is a common factor

OpenStudy (lgbasallote):

?

OpenStudy (zugzwang):

Scratch that proof, I think it's inaccurate.

OpenStudy (lgbasallote):

maybe \(\sqrt 3\) isn't irrational

OpenStudy (zugzwang):

lol it is We were pretty much on the right track... Ok, you've ruled out the possibility that they're both even. What about the other cases?

OpenStudy (zugzwang):

By both, I mean both a and b.

OpenStudy (zugzwang):

Suppose a is odd and b is even. then a^2 is odd and b^2 is even 3b^2 is even, but it is equal to a^2, an odd number, so this cannot be. Suppose a is even and b is odd Then a^2 is even and b^2 is odd 3b^2 is odd, but it is equal to a^2, an even number, so this cannot be. Hence we are left with the case that both a and b must be odd.

OpenStudy (zugzwang):

I realise you did that earlier... oh well...

OpenStudy (zugzwang):

Sorry for causing delay ... >.>

OpenStudy (lgbasallote):

hmm what about this a and b are both odd so... \[3(2k+1)^2 = (2x+1)^2\] \[3(4k^2 + 4k + 1) = 4x^2 + 4x + 1\] \[12k^2 + 12k + 3 = 4x^2 + 4x + 1\] \[12k^2 + 12k - 4x^2 + 4x = -2\] \[4(3k^2 + 3k - x^2 + x) = -2\] \[3k^2 + 3k - x^2 + x = -\frac 12\] contradiction

OpenStudy (zugzwang):

Something like that... Except in the third line, it should be -4x. Just make sure to indicate exactly why it must be a contradiction, and you're done.

OpenStudy (zugzwang):

Sorry, fourth line, not third.

OpenStudy (lgbasallote):

still contradicts though

OpenStudy (lgbasallote):

since you ahve the most repy in this thread, i'll medal you

OpenStudy (zugzwang):

lol, thanks :)